Europium metal crystallizes in a body–centered lattice with the unit cell measuring 4.58 

A° along each edge. Determine the volume of a unit cell.


1
Expert's answer
2021-06-07T03:15:27-0400

A body-centred lattice represents a cubic unit cell. According to the formula, the volume of a cube:

V = s3

where s is the edge length of the cube. 

From here, the volume of a unit cell equals:

V = s3 = (4.58 A°)3 = (4.58 × 10-10)3 = 9.607 × 10-29 m3


Answer: 9.607 × 10-29 m3